Solve 4 / 5 + 2 / 3 Solve 4 / 5 + 2 / 3 Solve 4 / 5 + 2 / 3

Solve
4 / 5 + 2 / 3
Solve
4 / 5 + 2 / 3
Solve
4 / 5 + 2 / 3

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

4 / 5 + 2 / 3
Taking LCM 15
= [(4 × 3) + (2 × 5)] / 15
= (12 + 10) / 15
= 22 / 15

Thanks
Shraddha Singh
3 years ago
Dear Student,
 
4/5 + 2/3 
Now we will take L.C.M.
L.C.M. of 3 and 5 is 15,
 
= 3×4 + 2×5 / 15
= 12 + 10 / 15
= 22/15